OpenStudy (anonymous):

What could be a real life example of basic related equations? REALLY URGENT!!

OpenStudy (anonymous):

like maybe an equation to triple or half a recipe?

OpenStudy (anonymous):

didnt think of that thanks!

OpenStudy (anonymous):

wait actually they have to be like writing out a bunch of different versions of one equation

OpenStudy (anonymous):

can you give me an example

OpenStudy (anonymous):

kk one sec

OpenStudy (anonymous):

x + 4=17 X=17-4 4=17-x The sum is 17. Subtract 4 from 17 for One equation, or you could subtract x from 17 for the other equation. x + 4=17 X=17-4 4=17-x The sum is 17. Subtract 4 from 17 for One equation, or you could subtract x from 17 for the other equation.

OpenStudy (anonymous):

sorry accidentally posted it twice

OpenStudy (anonymous):

hmm.... well for to half a recipe you can put x divided by 2=y y-x=? i tried but idk sorry :(

OpenStudy (anonymous):

thx though!